OpenStudy (anonymous):

Write the standard equation for the circle that passes through the points (1, 1), (1, 3), and (9, 2).

OpenStudy (anonymous):

use the formula of x^2 + y^2 + 2gx + 2fy + c = 0

OpenStudy (anonymous):

x 2 + y 2 + dx + ey + f = 0 this is the formula i have to use

OpenStudy (anonymous):

oh we were taught to use a diff formula but i suppose it still the same u get your first point and label it x and y then u sub the point into the equation so your first one would look like this (1)^2 + (1)^2 + d(1) + e(1) + f = 0

OpenStudy (anonymous):

ok i understand that much but I don't understand when it gets to the point of putting the answers of e and d and f back into the problem

OpenStudy (anonymous):

so after you have done that with each point u get three equations...then its basically just like three simultaneous equations and you get the answer by process of elimination

OpenStudy (anonymous):

then once you have e, d ans f u just sub it in to your original equation of x^2 + y^2 + dx + ey + f = 0